# An algorithm for estimating the crossing number of dense graphs, and continuous analogs of the crossing and rectilinear crossing numbers

### v2

@@ 3 Estimating the crossing number
A long edge of $G_{2}$ will be called lonely if it belongs to $E_{G_{2}}(V_{i},V_{j})$ and $e_{G_{1}}(V_{i},V_{j})\leq dn^{2}$. Note that $G_{2}$ has no more than $dk^{2}n^{2}\leq C^{2}d^{1/2}n^{2}$ lonely edges. Delete all the lonely edges from $\mathcal{D}^{\prime}$, we will later describe how they can be reinserted without increasing the crossing number by too much. A crossing point in either of $\mathcal{D}$ or $\mathcal{D}^{\prime}$ will be called bad it there is an $V_{i}$ that contains an endpoint of each of the two corresponding edges, and bad otherwise. There are three types of good crossings in $\mathcal{D}^{\prime}$: the ones that involve two non extremal sections (type $1$), those which involve a non extremal section and the interior of an extremal one (type $2$), and those that involve a short edge and a non extremal section (type $3$). Below, we bound expected value of $\mathop{\mathrm{cr}}(G_{2},\mathcal{D}^{\prime})$.
Every good crossing of type $1$ in $\mathcal{D}^{\prime}$ can be traced back to a good crossing in $\mathcal{D}$. If we think about it in this way, then a good crossing in $\mathcal{D}$ between an edge $(u,v)\in E_{G_{1}}(V_{i},V_{j})$ and an edge $(u^{\prime},v^{\prime})\in E_{G_{1}}(V_{s},V_{t})$ contributes at most
@@
to $\mathbf{E}[\mathop{\mathrm{cr}}(G_{2},\mathcal{D}^{\prime})]$. Let $c_{\mathcal{D}}(i,j,s,t)$ denote the total weighted sum of the crossings between an edge of $E_{G_{1}}(V_{i},V_{j})$ and an edge from $E_{G_{1}}(V_{s},V_{t})$ in $\mathcal{D}$. Then, the total contribution of these crossings to $\mathbf{E}[\mathop{\mathrm{cr}}(G_{2},\mathcal{D}^{\prime})]$ is upper bounded by
$c_{\mathcal{D}}(i,j,s,t)\frac{e_{G_{2}}(V_{i},V_{j})\cdot e_{G_{2}}(V_{s},V_{t})}{e_{G_{1}}(V_{i},V_{j})\cdot e_{G_{1}}(V_{s},V_{t})}\leq c_{\mathcal{D}}(i,j,s,t)\frac{(e_{G_{1}}(V_{i},V_{j})+dn^{2})(e_{G_{1}}(V_{s},V_{t})+dn^{2})}{e_{G_{1}}(V_{i},V_{j})\cdot e_{G_{1}}(V_{s},V_{t})}.$
Adding over all $4$-tuples $(i,j,s,t)$ with pairwise different entries (two tuples are considered the same if the correspond to the same pairs of parts of $\mathcal{P}$), we get that the expected weighted sum of the crossings of type 1 in $\mathcal{D}^{\prime}$ does not exceed
$\sum_{(i,j,s,t)}c_{\mathcal{D}}(i,j,s,t)\left(1+\frac{e_{G_{1}}(V_{i},V_{j})dn^{2}+e_{G_{1}}(V_{s},V_{t})dn^{2}+d^{2}n^{4}}{e_{G_{1}}(V_{i},V_{j})\cdot e_{G_{1}}(V_{s},V_{t})}\right)$
$\leq\sum_{(i,j,s,t)}c_{\mathcal{D}}(i,j,s,t)+e_{G_{1}}(V_{i},V_{j})dn^{2}+e_{G_{1}}(V_{s},V_{t})dn^{2}+d^{2}n^{4}$
@@
Next, we bound the amount of crossings of types $2$ and $3$ that are contained in $r_{i}$. Note that each of them can be traced back to an intersection point between a long edge of $G_{1}$ that has no endpoint in $V_{i}$ and a either a short edge of $G_{2}$ with endpoints in $V_{i}$ or an extremal section with endpoint in $V_{i}$ (i.e., an intersection point between an edge of $G_{1}$ with no endpoint in $V_{i}$ and an element of $E_{i}$). By construction, and thanks to property $3$ in the statement of Theorem 3.1, we have that, on average, each of the at most $\lceil\varepsilon^{2}n\rceil n$ elements of $E_{i}$ takes part in no more than $\varepsilon n^{2}$ crossings with the long edges of $G_{1}$ that have no endpoint in $V_{i}$. On top of this, if the long edge of $G_{1}$ involved in one of these crossings belongs to $E_{G_{1}}(V_{s},V_{t})$, then the crossing contributes at most
@@
to $\mathbf{E}[\mathop{\mathrm{cr}}(G_{2},\mathcal{D}^{\prime})]$. Since the lonely edges were deleted, we may assume that $e_{G_{1}}(V_{s},V_{t})>dn^{2}$, and thus the above quantity is less than $2$. Whence, the expected value of the weighted sum of the crossings of types $1$ and $2$ is less than $4\varepsilon^{3}n^{4}=4d^{3/8}n^{4}$ (recall that every edge has weight at most $1$).
Lastly, as any two edges with an endpoint in the same $V_{i}$ induce at most three crossings, the number of bad crossings does not surpass
$3k\binom{\lceil\varepsilon^{2}n\rceil}{2}\binom{n}{2}\leq 3k\varepsilon^{4}n^{4}\leq 3Cd^{1/4}n^{4}.$
Now we add the at most $C^{2}d^{1/2}n^{2}$ lonely edges back into the the drawing. In order to do this, we first transform $\mathcal{D}^{\prime}$ into a simple and locally optimal drawing by applying a sequence of operations that does not increase the weighted sum of the crossings. Then, each lonely edge can be represented by a curve that does not cross any non-lonely edge more than once, and we can get rid of multiple crossings between the lonely edges without increasing the number of crossings between lonely and non-lonely edges. At the end, we get a drawing $\mathcal{D}^{\prime\prime}$ of $G_{2}$ that satisfies
$\mathbf{E}[\mathop{\mathrm{cr}}(G_{2},\mathcal{D}^{\prime\prime}]\leq\mathop{\mathrm{cr}}(G_{2},\mathcal{D}^{\prime})+C^{2}d^{1/2}n^{2}\binom{n}{2}\leq\mathop{\mathrm{cr}}(G_{2},\mathcal{D}^{\prime})+C^{2}d^{1/2}n^{4}$
$\leq\mathop{\mathrm{cr}}(G_{1},\mathcal{D})+2C^{4}d^{1/2}n^{4}+4d^{3/8}n^{4}+3Cd^{1/4}n^{4}+C^{2}d^{1/2}n^{4}$
@@
where $M$ is an absolute constant. As this can be repeated for every drawing $\mathcal{D}$ of $G_{1}$, we arrive at $\mathop{\mathrm{cr}}(G_{2})\leq\mathop{\mathrm{cr}}(G_{1})+Md^{1/4}n^{4}$. Analogously, $\mathop{\mathrm{cr}}(G_{2})\leq\mathop{\mathrm{cr}}(G_{1})+Md^{1/4}n^{4}$. This proves the first part of the statement. For the second part, a direct computation yields $\mathop{\mathrm{cr}}(G_{2},\mathcal{D})\leq\mathop{\mathrm{cr}}(G_{1},\mathcal{D})+d^{2}n^{8}$, and we can finish as above. ∎
@@
For any edge weighted graph $G=(V,E)$ and any positive integer $m$, we have that
@@
This was shown to be true for the rectilinear crossing number (instead of the crossing number) in [18], the proof of this version is almost identical to the one that appears in the said paper.
We begin by proving the inequality on the left. Consider a simple drawing $\mathcal{D}$ of $G[m]$ which attains $\mathop{\mathrm{cr}}(G)$. From this drawing, we obtain a drawing $\mathcal{D}^{\prime}$ of $G$ in the following way: For each $v\in V$, choose uniformly at random an element $v^{\prime}$ from $U_{v}$ (see the definition of $G[m]$ in subsection 2.2); the point $p_{v}$ that represents $v^{\prime}$ in $\mathcal{D}$ will also be used to represent $v$ in $\mathcal{D}^{\prime}$. Once all these points have been chosen, notice that for every pair of adjacent vertices $u$ and $v$ of $G$ the points $p_{v}$ and $p_{u}$ are connected by a curve in $\mathcal{D}$; this is the curve that will represent the edge $(u,v)$ in $\mathcal{D}^{\prime}$. Clearly, $\mathcal{D}^{\prime}$ is simple. For any two non-adjacent edges $e_{1}=(a,b),e_{2}=(c,d)\in E$, the probability that they cross in $\mathcal{D}^{\prime}$ is $c(U_{a},U_{b},U_{c},U_{d})/m^{4}$, where $c(U_{a},U_{b},U_{c},U_{d})$ denotes the number of quadruples $(a^{\prime},b^{\prime},c^{\prime},d^{\prime})$ with $a^{\prime}\in U_{a}$, $b^{\prime}\in U_{b}$, $c^{\prime}\in U_{c}$, $d^{\prime}\in U_{d}$ such that the edges $(a^{\prime},b^{\prime})$ and $(c^{\prime},d^{\prime})$ cross each other in $\mathcal{D}$. From this, we get that
$\mathbf{E}[\mathop{\mathrm{cr}}(G,\mathcal{D}^{\prime})]\leq\mathop{\mathrm{cr}}(G[m],\mathcal{D})/m^{4}=\mathop{\mathrm{cr}}(G[m])/m^{4}.$
Since $\mathop{\mathrm{cr}}(G)\leq\mathbf{E}[\mathop{\mathrm{cr}}(G,\mathcal{D}^{\prime})]$, the result follows.
Now we prove the inequality on the right. Take a drawing $\mathcal{D}$ of $G$ that attains $\mathop{\mathrm{cr}}(G)$ and let $\varepsilon$ be a very small positive real (this will be made more precise in a moment). We obtain a drawing $\mathcal{D}^{\prime}$ of $G[m]$ as follows: For every $v\in V$, replace the point that represents $v$ by a set of $m$ distinct points, each of them at distance at most $\varepsilon$ from the original one; these points will represent the elements of $U_{v}$. For each $(u,v)\in E$, we draw all the edges between $U_{u}$ and $U_{v}$ by following along the curve that represents $(u,v)$ in $\mathcal{D}$ (this makes sense if $\varepsilon$ is small enough), making sure that no two edges of $G_{[}m]$ cross each other more than once. We call a $4$-tuple $(a,b,c,d)$ of vertices of $G[m]$ bad if two of its elements belong to the same $U_{v}$, and good otherwise. If $\varepsilon$ is small enough, then for any good $4$-tuple $(a,b,c,d)$ with $a\in U_{a}^{\prime}$, $b\in U_{b}^{\prime}$, $c\in U_{c}^{\prime}$, $d\in U_{d}^{\prime}$ the edges $(a,b)$ and $(c,d)$ cross in $\mathcal{D}^{\prime}$ if and only if $(a,b)$ and $(c,d)$ cross in $\mathcal{D}$. On the other hand, the number of bad $4$-tuples is less than $n\binom{m}{2}\binom{nm}{2}$, so we have at
@@
By theorems 3.3 and 3.4, the crossing density of graphs satisfies properties (i) and (ii) in subsection 2.3, and it definitely satisfies (iii) as well. This implies that the crossing density is an estimable graph parameter. That the rectilinear crossing density is estimable was already clear to the authors of [18].
@@
The algorithm
Here, we present the proof of Theorem 1.2. The algorithm we are about to describe is very similar to the one in [18] (see Theorem 1.1).
@@
    Consider the edge weighted graph $G/\mathcal{P}$ and compute its crossing number using Theorem 2.1. Then, output $cr(G/\mathcal{P})(n/k)^{4}$. This can be done in $2^{O(k^{4}\log k)}=n^{o(1)}$ time.
@@
    By Theorem 3.4, the blow-up $(G/\mathcal{P})[\lfloor n/K\rfloor]$ has crossing number at most $\text{cr}(G/\mathcal{P})(n/k)^{4}$. The graph $(G/\mathcal{P})[\lfloor n/K\rfloor]$ can be obtained from $G_{\mathcal{P}}$ by removing less than $k$ vertices (along with the attached edges), thus $\text{cr}(G_{\mathcal{P}})\leq(G/\mathcal{P})[\lfloor n/K\rfloor]+kn^{3}$. Finally, $\mathcal{P}$ was chosen so that $|G-G_{\mathcal{P}}|\leq\varepsilon$, so Theorem 3.3 yields $|\text{cr}(G)-\text{cr}(G_{\mathcal{P}})|\leq M\varepsilon^{1/4}n^{4}$. These observations imply that
    $\text{cr}(G)\leq\text{cr}(G/\mathcal{P})(n/k)^{4}+kn^{3}+M\varepsilon^{1/4}n^{4}\leq\text{cr}(G/\mathcal{P})(n/k)^{4}+O\left(\frac{n^{4}}{(\log\log n)^{\frac{1}{8c}}}\right).$
@@
    If we also wish to find a drawing of $G$ with $\mathop{\mathrm{cr}}(G)+O(n^{4}/(\log\log n)^{\delta^{\prime}})$ crossings, then we first need access to a simple and locally optimal drawing of drawing of $G_{\mathcal{P}}$ (or a graph that is very close to $G_{\mathcal{P}}$ with respect to $d_{\square}$) with crossing number $\mathop{\mathrm{cr}}(G_{\mathcal{P}})+O(n^{4}/(\log\log n)^{\delta^{\prime}})$. A drawing of $G_{\mathcal{P}}$ with this many crossings can be obtained by combining the drawing of $G/\mathcal{P}$ provided by 2.1 with the technique used in the proof of the right side inequality in the statement of Theorem 3.4 (see the above proof). Now, we round the weights of the edges so that they can be written as an integer divided by $O(1/(\log\log n)^{\delta^{\prime}/2})$, thus obtaining a new graph $G_{\mathcal{P}}^{\prime}$ whose crossing number differs from that of $G_{\mathcal{P}}$ by at most $O(n^{4}/(\log\log n)^{\delta^{\prime}})$. In order to construct a locally optimal drawing of $G_{\mathcal{P}}^{\prime}$, we iteratively refine the drawing by using Dijkstra’s algorithm to redraw the edges until it is no longer possible to keep going. Each step reduces the weighted sum of the crossings by $\Omega(1/(\log\log n)^{\delta^{\prime}})$, so the whole process takes polynomial-time. Now we can follow the proof of Theorem 3.3, which is essentially algorithmic, to obtain a drawing of $G$. At some point during this procedure, we also require a locally optimal drawing of a certain subgraph of $G$; this drawing can be constructed just as we did for $G_{\mathcal{P}}^{\prime}$, since every step reduces the weighted sum of the crossings by at least $1$. The expected number of crossings in the resulting drawing of $G$ is $\mathop{\mathrm{cr}}(G)+o(n^{4})$, and the probability that the algorithm fails to deliver a drawing with crossing number close to this quantity can be bounded from above using the fact that the no drawing with crossing number less than $\mathop{\mathrm{cr}}(G)$ can ever be generated.
    Remark: The only step of the above algorithm at which we produce an actual graph drawing is at the end. Before that, drawings are processed not as geometric objects, but combinatorial ones (in the sense that they are stored as planar maps with a vertex for each crossing point and for each vertex of the original graph). The final drawing can be constructed using any efficient algorithm for drawing planar graphs (see Chapter 6 in [41] for many such algorithms).

### v4

@@ 3.2 Cut distance and crossing numbers
The goal now is to bound the expected value of $\mathop{\mathrm{cr}}(G_{2},\mathcal{D}^{\prime})$. A long edge of $G_{2}$ will be called lonely if it belongs to $E_{G_{2}}(V_{i},V_{j})$ and $e_{G_{1}}(V_{i},V_{j})\leq dn^{2}$. Note that $G_{2}$ has no more than $dk^{2}n^{2}\leq C^{2}d^{1/2}n^{2}$ lonely edges. Delete all the lonely edges from $\mathcal{D}^{\prime}$; we will later describe how they can be reinserted without creating too many crossings. There are four types of crossings in $\mathcal{D}^{\prime}$: the ones that involve two non-extremal sections (type $1$), those which involve a non-extremal section and the interior of an extremal one (type $2$), those that involve a short edge and a non-extremal section (type $3$), and finally the ones that involve only short edges and extremal sections (type 4).
Every crossing of type $1$ in $\mathcal{D}^{\prime}$ can be traced back to a crossing in $\mathcal{D}$. If we think about it in this way, then a crossing in $\mathcal{D}$ between an edge $(u,v)\in E_{G_{1}}(V_{i},V_{j})$ and an edge $(u^{\prime},v^{\prime})\in E_{G_{1}}(V_{s},V_{t})$ contributes at most
@@
to $\mathbb{E}[\mathop{\mathrm{cr}}(G_{2},\mathcal{D}^{\prime})]$. Let $c_{\mathcal{D}}(i,j|s,t)$ denote the total weighted sum of the crossings between an edge of $E_{G_{1}}(V_{i},V_{j})$ and an edge from $E_{G_{1}}(V_{s},V_{t})$ in $\mathcal{D}$. Then, the total contribution of these crossings to $\mathbb{E}[\mathop{\mathrm{cr}}(G_{2},\mathcal{D}^{\prime})]$ is upper bounded by
$c_{\mathcal{D}}(i,j|s,t)\frac{e_{G_{2}}(V_{i},V_{j})\cdot e_{G_{2}}(V_{s},V_{t})}{e_{G_{1}}(V_{i},V_{j})\cdot e_{G_{1}}(V_{s},V_{t})}\leq c_{\mathcal{D}}(i,j|s,t)\frac{(e_{G_{1}}(V_{i},V_{j})+dn^{2})(e_{G_{1}}(V_{s},V_{t})+dn^{2})}{e_{G_{1}}(V_{i},V_{j})\cdot e_{G_{1}}(V_{s},V_{t})}.$
Adding over all $4$-tuples $(i,j,s,t)$ of numbers in $\{1,\dots,k\}$ (two tuples are considered the same if the correspond to the same pairs of parts of $\mathcal{P}$), we get that the expected weighted sum of the crossings of type 1 in $\mathcal{D}^{\prime}$ does not exceed
$\sum_{(i,j,s,t)}c_{\mathcal{D}}(i,j|s,t)\left(1+\frac{e_{G_{1}}(V_{i},V_{j})dn^{2}+e_{G_{1}}(V_{s},V_{t})dn^{2}+d^{2}n^{4}}{e_{G_{1}}(V_{i},V_{j})\cdot e_{G_{1}}(V_{s},V_{t})}\right)$
$\leq\sum_{(i,j,s,t)}c_{\mathcal{D}}(i,j|s,t)+e_{G_{1}}(V_{i},V_{j})dn^{2}+e_{G_{1}}(V_{s},V_{t})dn^{2}+d^{2}n^{4}$
@@
Next, we bound the amount of crossings of types $2$ and $3$ that are contained in $r_{i}$. Note that each of them can be traced back to an intersection point between a long edge of $G_{1}$ that has no endpoint in $V_{i}$ and a either a short edge of $G_{2}$ with endpoints in $V_{i}$ or an extremal section with endpoint in $V_{i}$ (in other words, an intersection point between an edge of $G_{1}$ with no endpoint in $V_{i}$ and an element of $E_{i}$). By construction, and thanks to property (III) in the statement of Theorem 3.1, we have that on average each of the at most $\lceil\varepsilon^{2}n\rceil n$ elements of $E_{i}$ takes part in no more than $\varepsilon n^{2}$ crossings with the long edges of $G_{1}$ that have no endpoint in $V_{i}$. On top of this, if the long edge of $G_{1}$ involved in one of these crossings belongs to $E_{G_{1}}(V_{s},V_{t})$, then the crossing contributes at most
@@
to $\mathbb{E}[\mathop{\mathrm{cr}}(G_{2},\mathcal{D}^{\prime})]$. Since the lonely edges were deleted, we may assume that $e_{G_{1}}(V_{s},V_{t})>dn^{2}$, and thus the above quantity is less than $2$. Whence, the expected value of the weighted sum of the crossings of types $2$ and $3$ is upper bounded by $2\lceil\varepsilon^{2}n\rceil n\cdot\varepsilon n^{2}<4\varepsilon^{3}n^{4}=4d^{3/8}n^{4}$.
Lastly, as any two short edges or extremal sections that are fully contained in $r_{i}$ induce at most $1$ crossing, the number of crossings of type $4$ does not surpass
$k\binom{\lceil\varepsilon^{2}n\rceil}{2}\binom{n}{2}\leq k\varepsilon^{4}n^{4}\leq Cd^{1/4}n^{4}.$
Now, we add the at most $C^{2}d^{1/2}n^{2}$ lonely edges back into the drawing. In order to do this, we first transform $\mathcal{D}^{\prime}$ into a simple and locally optimal drawing by applying a sequence of operations that does not increase the weighted sum of the crossings. At this point, each lonely edge can be represented by a curve that does not cross any non-lonely edge more than once. We can get rid of multiple crossings between the lonely edges by rerouting some of the curves without increasing the number of crossings between lonely and non-lonely edges. At the end, we get a drawing $\mathcal{D}^{\prime\prime}$ of $G_{2}$ that satisfies
$\mathbb{E}[\mathop{\mathrm{cr}}(G_{2},\mathcal{D}^{\prime\prime}]\leq\mathbb{E}[\mathop{\mathrm{cr}}(G_{2},\mathcal{D}^{\prime})]+C^{2}d^{1/2}n^{2}\binom{n}{2}\leq\mathbb{E}[\mathop{\mathrm{cr}}(G_{2},\mathcal{D}^{\prime})]+C^{2}d^{1/2}n^{4}$
$\leq\mathop{\mathrm{cr}}(G_{1},\mathcal{D})+2C^{4}d^{1/2}n^{4}+4d^{3/8}n^{4}+Cd^{1/4}n^{4}+C^{2}d^{1/2}n^{4}$
@@
where $M$ is an absolute constant. As this can be repeated for every drawing $\mathcal{D}$ of $G_{1}$, we arrive at $\mathop{\mathrm{cr}}(G_{2})\leq\mathop{\mathrm{cr}}(G_{1})+Md^{1/4}n^{4}$. Analogously, $\mathop{\mathrm{cr}}(G_{2})\leq\mathop{\mathrm{cr}}(G_{1})+Md^{1/4}n^{4}$. This proves the first part of the statement. For the second part, a direct computation yields $\mathop{\mathrm{cr}}(G_{2},\mathcal{D})\leq\mathop{\mathrm{cr}}(G_{1},\mathcal{D})+d^{2}n^{8}$ (that is, we do not need to modify the drawing of $G_{1}$ at all to get the desired inequality), and we can finish as above . ∎
Next, we address how taking the blow-up of a graph affects the crossing number.
@@
For any edge-weighted $n$-vertex graph $G=(V,E)$ and any positive integer $m$, we have that
@@
Remark. This was shown to be true for the rectilinear crossing number (instead of the crossing number) in [21]; the proofs are very similar, and the fact that this holds also for the crossing number was already known to Fox, Pach and Suk.
@@
We begin by proving the inequality on the left. Consider a simple drawing $\mathcal{D}$ of $G[m]$ which attains $\mathop{\mathrm{cr}}(G)$. From this drawing, we obtain a drawing $\mathcal{D}^{\prime}$ of $G$ in the following way: Independently for each $v\in V$, choose uniformly at random an element $v^{\prime}$ from $U_{v}$ (see the definition of $G[m]$ in Section 2.2); the point $p_{v}$ that represents $v^{\prime}$ in $\mathcal{D}$ will be used to represent $v$ in $\mathcal{D}^{\prime}$. Once all these points have been chosen, notice that for every two adjacent vertices $u$ and $v$ of $G$ the points $p_{v}$ and $p_{u}$ are connected by a curve in $\mathcal{D}$; this is the curve that will represent the edge $(u,v)$ in $\mathcal{D}^{\prime}$. Clearly, $\mathcal{D}^{\prime}$ is simple. For any two non-adjacent edges $e_{1}=(a,b),e_{2}=(c,d)\in E$, the probability that they cross in $\mathcal{D}^{\prime}$ is $c(U_{a},U_{b}|U_{c},U_{d})/m^{4}$, where $c(U_{a},U_{b}|U_{c},U_{d})$ denotes the number of quadruples $(a^{\prime},b^{\prime},c^{\prime},d^{\prime})$ with $a^{\prime}\in U_{a}$, $b^{\prime}\in U_{b}$, $c^{\prime}\in U_{c}$, $d^{\prime}\in U_{d}$ such that the edges $(a^{\prime},b^{\prime})$ and $(c^{\prime},d^{\prime})$ cross each other in $\mathcal{D}$. From this, we get that
$\mathbb{E}[\mathop{\mathrm{cr}}(G,\mathcal{D}^{\prime})]\leq\mathop{\mathrm{cr}}(G[m],\mathcal{D})/m^{4}=\mathop{\mathrm{cr}}(G[m])/m^{4}.$
Since $\mathop{\mathrm{cr}}(G)\leq\mathbb{E}[\mathop{\mathrm{cr}}(G,\mathcal{D}^{\prime})]$, the result follows.
Now we prove the inequality on the right. Take a drawing $\mathcal{D}$ of $G$ that attains $\mathop{\mathrm{cr}}(G)$ and let $\varepsilon$ be a very small positive real (this will be made more precise in a moment). We obtain a drawing $\mathcal{D}^{\prime}$ of $G[m]$ as follows: For every $v\in V$, replace the point that represents $v$ by a set of $m$ distinct points, each of them at distance at most $\varepsilon$ from the original one; these points will represent the elements of $U_{v}$. For each $(u,v)\in E$, we draw all the edges between $U_{u}$ and $U_{v}$ by following along the curve that represents $(u,v)$ in $\mathcal{D}$ (this makes sense if $\varepsilon$ is small enough), making sure that no two edges of $G[m]$ cross each other more than once. We call a $4$-tuple $(a,b,c,d)$ of vertices of $G[m]$ bad if two of its elements belong to the same $U_{v}$, and good otherwise. If $\varepsilon$ is small enough, then for any good $4$-tuple $(a,b,c,d)$ with $a\in U_{a}^{\prime}$, $b\in U_{b}^{\prime}$, $c\in U_{c}^{\prime}$, $d\in U_{d}^{\prime}$ the edges $(a,b)$ and $(c,d)$ cross in $\mathcal{D}^{\prime}$ if and only if $(a,b)$ and $(c,d)$ cross in $\mathcal{D}$. On the other hand, the number of crossings coming from bad $4$-tuples is less than $n\binom{m}{2}\binom{nm}{2}$, so we have that
@@
By theorems 3.3 and 3.4, the crossing density of graphs satisfies properties (i) and (ii) in Section 2.3, and it clearly satisfies (iii) as well. This implies that the crossing density is an estimable graph parameter. That the rectilinear crossing density is estimable was already shown in [21].
@@
3.3 The algorithm
Here, we present the algorithm that was promised in Theorem 1.2. The first three steps of the algorithm are completely standard and, in particular, very similar to the algorithm from [21] (see Theorem 1.1).
@@ 3.3 The algorithm
    Consider the edge-weighted graph $G/\mathcal{P}$ and compute its crossing number using Theorem 2.1. Then, output $\text{cr}(G/\mathcal{P})\cdot(\lfloor n/k\rfloor)^{4}$. This can be done in $2^{O(k^{4}\log k)}=n^{o(1)}$ time.
@@
    Let $H$ denote the blow-up graph $(G/\mathcal{P})[\lfloor n/k\rfloor]$. By Theorem 3.4, the crossing number of $H$ satisfies:
    $\text{cr}(G/\mathcal{P})(\lfloor n/k\rfloor)^{4}\leq\text{cr}(H)\leq\text{cr}(G/\mathcal{P})(\lfloor n/k\rfloor)^{4}+n^{4}/k.$
    The graph $H$ can be obtained from $G_{\mathcal{P}}$ by removing less than $k$ vertices (along with the attached edges), thus
    $\text{cr}(H)\leq\text{cr}(G_{\mathcal{P}})\leq\text{cr}(H)+kn^{3}.$
    Finally, $\mathcal{P}$ was chosen so that $d_{\square}(G,G_{\mathcal{P}})\leq\varepsilon$, so Theorem 3.3 yields $|\text{cr}(G)-\text{cr}(G_{\mathcal{P}})|\leq M\varepsilon^{1/4}n^{4}$. Together, these inequalities imply that
    $|\text{cr}(G/\mathcal{P})(\lfloor n/k\rfloor)^{4}-\text{cr}(G)|\leq n^{4}/k+kn^{3}+M\varepsilon^{1/4}n^{4}=O\left(\frac{n^{4}}{(\log\log n)^{\frac{1}{8c}}}\right).$
@@
    Write $\delta^{\prime}=1/8c$. If we also wish to find a drawing of $G$ with $\mathop{\mathrm{cr}}(G)+O(n^{4}/(\log\log n)^{\delta^{\prime}})$ crossings, then we first need access to a simple and locally optimal drawing of $G_{\mathcal{P}}$ (or of a graph that is very close to $G_{\mathcal{P}}$ with respect to $d_{\square}$) with crossing number $\mathop{\mathrm{cr}}(G_{\mathcal{P}})+O(n^{4}/(\log\log n)^{\delta^{\prime}})$. A drawing of $G_{\mathcal{P}}$ with this many crossings can be obtained by combining the drawing of $G/\mathcal{P}$ provided by 2.1 with the technique used in the proof of the right hand side inequality in the statement of Theorem 3.4. Next, we round the weights of the edges of $G_{\mathcal{P}}$ so that they can be written as an integer divided by some fixed integer $s=\Theta(\log\log n)^{\delta^{\prime}}$, thus obtaining a new graph $G_{\mathcal{P}}^{\prime}$ whose crossing number differs from that of $G_{\mathcal{P}}$ by at most $O(n^{4}/(\log\log n)^{\delta^{\prime}})$. In order to construct a locally optimal drawing of $G_{\mathcal{P}}^{\prime}$, we iteratively refine the drawing by using Dijkstra’s algorithm to redraw the edges until it is no longer possible to reduce the crossing number of the drawing in this manner. Each step reduces the weighted sum of the crossings by at least $1/s^{2}$, so the whole process takes polynomial time. Now, we can follow the proof of Theorem 3.3, which is essentially algorithmic, to obtain a random drawing of $G$. At some point during this procedure, we also require a locally optimal drawing of a certain subgraph of $G$; this drawing can be constructed just as we did for $G_{\mathcal{P}}^{\prime}$, since every step will reduce the weighted sum of the crossings by at least $1$. The expected number of crossings in the resulting drawing of $G$ is $\mathop{\mathrm{cr}}(G)+O(n^{4}/(\log\log n)^{\delta^{\prime}})$, and the probability that the algorithm fails to deliver a drawing with crossing number close to this quantity can be bounded from above using Markov’s inequality together with the fact that no drawing with crossing number less than $\mathop{\mathrm{cr}}(G)$ can ever be generated.
    Remark. The only step in part $4$ of the above algorithm at which we produce an actual graph drawing is at the end. Before that, drawings are processed not as geometric objects, but combinatorial ones (in the sense that they are stored as planar maps with a vertex for each crossing point and for each vertex of the original graph). The final drawing can be constructed using any efficient algorithm for drawing planar graphs (see Chapter 6 in [49] for many such algorithms).